1)  Computing labor force statistics:

Adult population of the U.S. by group, 2019

# of employed 151.1 million

# of unemployed 7.8 million

not in labor force 94.5 million

1. Based on the above table, compute the labor force, u-rate, adult population, and labor force participation rate. Evaluating “U” measure:

2. In each of the following, what happens to the u-rate? Does the u-rate give an accurate impression of what’s happening in the labor market? 

A. Sue lost her job and begins looking for a new one.

B. Jon, a steelworker who has been out of work since his mill closed last year, becomes discouraged and gives up looking for work.

C. Sam, the sole earner in his family of 5, just lost his $80,000 job as a research scientist. Immediately, he takes a part-time job at McDonald’s until he can find another job in his field. Natural unemployment:

 



1
Expert's answer
2021-09-19T18:35:15-0400

1. The labor force is: LF = E + U = 151.1 + 7.8 = 158.9 million.

The u-rate is:

u = 7.8/158.9×100% = 4.9%.

Adult population is:

158.9 + 94.5 = 253.4 million.

And labor force participation rate is:

LFPR = 158.9/253.4×100% = 62.71%.

2. A) u-rate will increase.

B) u-rate will decrease.

C) no change.